OpenStudy (anonymous):

mike and steve have wallets with five and ten dollar bills in them. If together, they have 20 bills total and it tatals $135, how many $5's and 10's do they have?

OpenStudy (campbell_st):

you have to solve 2 equations simultaneously let f = $5 and t = $10 f + t = 20 number of bills 5f + 10t = 135 total value easiest method is to rewrite the 1st equation t = 20 - f and substitute it into the 2nd equation 5f + 10(20 - f) = 135 distribute and solve for f when you have the number of $5 then substitute that number into the 1st equation to get the number of $10. hope this helps

OpenStudy (anonymous):

Distribute?

OpenStudy (anonymous):

distribute what?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

the 10(20 - f) so the equation is 5f + 200 - 10f = 135 collect like terms and solve for f

OpenStudy (anonymous):

Oh, okay, thanks:)